SINGLET-TRIPLET SPECTROSCOPY OF BUTADIENE

\underline{PAUL R. WINTER}, ALLISON G. ROBINSON, TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

The T_{1}-S_{0} absorption spectrum of gas phase 1,3-butadiene was recorded over the region from 20,500 to 23,000 cm^{−1} using cavity ring-down spectroscopy. The origin is located at 20,777 cm^{−1}, and several vibronic bands are observed. These bands are assigned to vibrations of a planar triplet structure based on Franck- Condon calculations. This structure has been calculated to be about 3 kcal mol^{−1} higher in energy than the twisted structure, in which one CH_{2} group is oriented perpendicular to the plane of the rest of the molecule. A two-dimensional torsional calculation involving the rotation of the two terminal CH_{2} groups suggests that the planar structure is located in a shallow minimum at the top of a barrier between two twisted structures.